# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v3

@@ 2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis. These studies collectively form the foundation for our own framework SARA, which builds upon their methods but advances them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
@@
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling. Ghosh et al. (Ghosh_Pargaonkar_Eisty_2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al_Wahshat_Abu_ulbeh_Yusoff_Zakaria_Amir_Fazamin_Wan_Hamzah_P_2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
@@
Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment. While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation. Our filtering component achieves comparable accuracy (90%) while operating at a much larger scale across diverse Gen-AI app categories.
@@
Pre-LLM approaches to topic extraction. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi_Abran_2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (10499727) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modeling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi_Hassan_Tian_Zou_2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
@@
LLM-based topic extraction and assignment methods. LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola_Lalendra_Wimalasena_Munasinghe_2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham_Hoyle_Sun_Resnik_Iyyer_2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (openai2023gpt) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash_Wang_Hoang_Hee_Lee_2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews. Assi et al. (Maram_LLM_Cure) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Among these, LLM-Cure most closely resembles our framework in combining both extraction and assignment tasks, though its focus was on generating feature suggestions rather than large-scale trend analysis.
@@
Single-task topic analysis. Instead of extracting and assigning topics to user reviews, some studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei_Courbis_Lambolais_Xu_Bernard_Dray_2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023) and Ren et al. (Ren_Nakagawa_Tsuchiya_2024) extracted topics from user reviews without performing the assignment task. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
@@
Domain-specific applications of LLMs in user review analysis. Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis. Sorathiya and Ginde (Sorathiya_Ginde_2024) employed a hybrid approach combining Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss. Rezaei Nasab et al. (RezaeiNasab_Dashti_Shahin_Zahedi_Khalajzadeh_Arora_Liang_2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning based classifiers to detect fairness mentions and cluster them into six categories.
@@
While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially. We analyze user reviews from a large-scale collection of Gen-AI apps spanning multiple functional categories, including but not limited to health & fitness, education, productivity, and entertainment. Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e. AI performance, content quality, and content policy) that are unique to this new class of applications. This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
@@ 3. Methodology
To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning framework to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. SARA integrates and extends state-of-the-art methods in user review analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), combining automated filtering, topic extraction, and topic assignment into a unified, empirically validated workflow that supports large-scale, LLM-driven analysis of user feedback. Although evaluated within the Gen-AI domain, SARA’s modular design enables its application to other mobile or AI-enabled systems for analyzing large-scale user feedback. We first evaluate the reliability of SARA in accurately labeling review topics (RQ1) and then apply it in subsequent analyses to explore user concerns and temporal trends (RQ2–RQ3). We present an overview of our research framework in Figure 2.
@@ 3.1. Selection of Gen-AI apps
Our app selection process involves three steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, and 3) manually screening and validating each app to confirm it is Gen-AI.
@@
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (gozalobrizuela2023surveygenerativeaiapplications). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
    Figure 2. Overview of our framework
@@
    App scraping. Using our list of keywords, we scrape the Google Play Store to identify all matching apps. Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store. For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description. In this paper, we refer to Google Play Store app categories as app categories.
@@
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (cohen1960kappa) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
Category Keywords
Image Image Editing, Artistic Images, Realistic Images
Video Text-to-Video, Video Production, Image-to-Video
3D Text-to-3D
Code and Software Text-to-Code, Text-to-Multilingual Code
Speech Text-to-Speech, AI Understanding, Speech-to-Text
Business Business
Marketing Marketing, Advertisement, Accounting
Gaming Video game Creation
Music Music Generator
@@ 3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (googleplayscraper). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
@@ 3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning framework designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
@@
Table 2. Number of reviews by app category and stage (Full dataset considered).
App category Stage 0 Stage 1 Stage 2 Stage 3
Productivity 362,210 354,829 233,311 169,119
Photography 1,035,946 430,490 188,382 126,277
Art & Design 332,881 238,291 157,320 125,362
Entertainment 205,467 205,467 156,636 118,669
Education 101,812 95,590 57,680 41,905
Video Players & Editors 364,124 109,610 51,182 36,507
Tools 231,865 104,636 48,807 35,046
Music & Audio 17,875 17,875 11,781 9,150
Health & Fitness 41,537 10,368 8,747 7,849
Games 29,922 6,254 5,555 5,070
Books & Reference 2,071 2,071 1,423 1,112
Total 2,725,710 1,575,481 920,824 676,066
@@
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh_Pargaonkar_Eisty_2024; Prakash_Wang_Hoang_Hee_Lee_2023; Roumeliotis_Tselikas_Nasiopoulos_2024; Lee_Lee_Seo_2024), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labeled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
@@
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam_Ingio_2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar_Hasan_Kumar_Carroll_2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi_Elluri_2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar_2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
@@
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi_Hassan_Tian_Zou_2021; 10499727). This step aims to eliminate overly brief and potentially non-informative feedback.
@@
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen_Lin_Hoi_Xiao_Zhang_2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
@@
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (openai2023gpt) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI_O4Mini_2024).
    Prior research comparing explicit, implicit, and creative prompting (Giray_2023; Ho_Mayberry_Nguyen_Dhulipala_Pallipuram_2024; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning. Consistent with these findings and with approaches such as LLM-Cure (Maram_LLM_Cure), we design an explicit, structured prompt that classifies each review as either informative or non-informative. The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses. These measures ensure deterministic labeling and minimize hallucination by preventing free-form or extraneous text generation.
@@
    Iterative prompt refinement. Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; He_Rungta_Koleczek_Sekhon_Wang_Hasan_2024). In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (cohen1960kappa) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.

### v4

@@ 2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis. These studies collectively form the foundation for our own framework, SARA, which builds upon their methods while advancing them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
@@
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling. Ghosh et al. (Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
@@
Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment. While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation. Our filtering component achieves comparable accuracy (91%) while operating at a much larger scale across diverse Gen-AI app categories.
@@
Pre-LLM approaches to topic extraction. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a 2016 literature review exploring approaches to mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews as positive, neutral, or negative. They then used an LDA model for topic modeling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi et al., 2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79-82% precision and 74-77% recall on annotated review data.
@@
LLM-based topic extraction and assignment methods. LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews. Assi et al. (Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Among these, LLM-Cure most closely resembles our framework in combining both extraction and assignment tasks, though its focus was on generating feature suggestions rather than large-scale trend analysis.
@@
Single-task topic analysis. Instead of extracting and assigning topics to user reviews, some studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos Santos et al., 2023) and Ren et al. (Ren et al., 2024) extracted topics from user reviews without performing the assignment task. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
@@
Domain-specific applications of LLMs in user review analysis. Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis. Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combined Natural Language Inference (NLI) with LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss. Rezaei Nasab et al. (Rezaei Nasab et al., 2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning-based classifiers to detect fairness mentions and cluster them into six categories.
@@
While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially. We analyze user reviews from a large-scale collection of Gen-AI apps across multiple functional categories, including education, productivity, and entertainment. Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e. AI performance, content quality, and content policy) that are unique to this new class of apps. This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
@@ 3. Methodology
To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. We present an overview of our research framework in Figure 2.
@@ 3.1. Selection of Gen-AI apps
Our app selection process involves four steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, 3) manually screening and validating each app to confirm that it satisfies our definition of a Gen-AI app as introduced earlier, and 4) finding the App Store version of that app. Specifically, we retain only apps whose core functionality involves generating novel content (e.g., text, images, audio, or video) and exclude apps that rely solely on AI for prediction, analysis, or classification without generating new content.
Figure 2. Overview of our framework.
An overview diagram of the SARA framework. The process begins with the Google Play Store, where Gen-AI apps are selected through four steps: (1) generating keywords, (2) scraping Google Play Store apps, (3) screening and validating apps, and (4) identifying corresponding App Store apps. This selection process results in 194 confirmed Gen-AI apps. User reviews are then collected, yielding 4,312,168 reviews from 193 apps. These reviews are refined through a multi-stage filtering process, resulting in 1,035,342 informative reviews from 171 apps. Finally, the refined dataset is analyzed through three research questions: (RQ1) designing prompts and evaluating accuracy, (RQ2) identifying and analyzing Gen-AI topics, and (RQ3) examining the evolution of these topics over time.
@@
    Generate Keywords. We compile a list of keywords derived from a Gen-AI taxonomy proposed by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Merchan, 2024). We adapt this taxonomy to mobile app contexts by selecting relevant categories and deriving corresponding keywords. The complete list of categories and keywords is provided in Appendix A.
@@
    Scrape Google Play Store Apps. Using our list of keywords, we scrape the Google Play Store to identify all matching apps. For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description. In this paper, we refer to Google Play Store app categories as app categories. Our app data collection occurred in October 2024.
@@
    Screen and Validate Apps. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by examining their titles, descriptions, and images. We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve discrepancies in ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
  4. 4.
    Identify the Corresponding App Store Apps. Our empirical study primarily focuses on user reviews from the Google Play Store. However, we also identify the corresponding App Store versions of these apps in order to examine whether the identified themes generalize across platforms. Out of the 194 Gen-AI apps from the Google Play Store, we identify the App Store version using AppMagic  [^4] https://appmagic.rocks/top-charts/apps?tag=243926&store=1&date=2024-10-01&topDepth=1000. We identify 81 App Store apps in January 2026.
@@ 3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, n.d.). For each retrieved review, the crawler extracts a comprehensive set of metadata, including content (i.e., review text), score (i.e., user rating from 1 to 5 stars), at (i.e., timestamp of the review), replyContent (developer’s response), and repliedAt (i.e., timestamp of response). We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 4,312,168 reviews of 193 apps. User reviews from the App Store apps are collected using the Appbot platform (Appbot, 2026). We extract the same review metadata as Google Play Store reviews. At the end of this step, we collect 540,119 reviews from 70 apps. We map App Store reviews to Google Play Store app categories to enable later cross-platform contextualization. Review acquisition occurred in February 2026.
@@ 3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data-cleaning framework that progressively improves the quality of user review data. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 1 for Google Play Store reviews and in Table 2 for App Store reviews. Each stage builds upon the previous one as follows:
@@
Table 1. Number of Google Play Store reviews by app category and stage (Full dataset considered).
App category Processing Stages
Stage 0 Stage 1 Stage 2 Stage 3
Basic cleaning Temporal filtering short review exclusion LLM filtering
Productivity 903,719 817,968 599,382 402,055
Photography 1,349,480 613,680 284,065 170,446
Entertainment 431,740 257,237 205,108 152,139
Art & Design 509,505 302,899 174,911 125,375
Education 156,435 142,047 95,254 65,699
Video Players & Editors 476,730 136,542 62,807 39,972
Tools 271,218 94,374 52,070 36,694
Games 35,853 34,850 29,648 25,211
Music & Audio 48,667 44,487 25,781 17,751
Total 4,183,347 2,444,084 1,529,026 1,035,342
Table 2. Number of App Store reviews by app category and stage (Full dataset considered).
App category Processing Stages
Stage 0 Stage 1 Stage 2 Stage 3
Basic cleaning Temporal filtering short review exclusion LLM filtering
Productivity 147,288 136,807 122,580 78,767
Photography 109,752 26,639 23,028 15,375
Entertainment 60,939 37,987 36,522 26,777
Art & Design 74,311 51,700 48,062 33,556
Education 117,752 115,121 92,002 56,944
Video Players & Editors 9,763 7,608 6,848 4,420
Tools 4,329 3,826 3,604 2,504
Games 5,328 5,083 4,910 3,774
Music & Audio 7,317 7,016 5,930 3,352
Total 536,779 391,787 343,486 225,469
@@
Stage 0: Basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis. In the initial stage, following prior work, we normalize whitespace and Unicode characters; remove emojis, URLs, and usernames; replace commas with whitespace; convert to lowercase; and drop empty reviews.
@@
Stage 1: Temporal filtering. This stage filters reviews written prior to Gen-AI integration to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^5] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. We outline our temporal filtering procedure below.
  • Integration Date Identification. To approximate the Gen-AI integration date for each app, we adopt a conservative, multi-source identification strategy and record the earliest verifiable date across all available evidence. Specifically, we consult the following sources: (i) developer websites, blog posts, press releases, and public announcements; and (ii) historical release notes and version histories obtained from AppMagic [^6] https://appmagic.rocks/. When multiple sources report different dates, we retain the earliest date supported by documentary evidence. This approach prioritizes evidentiary certainty over speculative precision.
  • Fallback Strategy. For Google Play Store apps, if no explicit integration date can be identified, or if the earliest documented evidence postdates October 2024, we conservatively assign October 2024 as the reference date. This date corresponds to the time at which we independently screened and verified the presence of Gen-AI functionality in the app. Therefore, the recorded date reflects the earliest verifiable confirmation of Gen-AI features rather than the exact deployment date. For App Store apps, we do not apply the October 2024 fallback, as App Store descriptions were not archived at the time of app selection. Consequently, if no verifiable integration date could be identified using release notes or developer sources, the app is excluded from temporal analysis.
  • Summary of Identified Dates. For Google Play Store apps, we identify integration dates for 4 apps through public articles, for 125 apps through release notes, and 64 apps are assigned to October 2024 based on independent verification. For App Store apps, integration dates are identified for 2 apps via developer websites, 5 through public articles, and 63 through release notes. We exclude 10 App Store apps due to the absence of verifiable integration dates.
  • Reliability and Coding Procedure. To assess reliability, we randomly sample 65 apps from our 193-app list (95% confidence level, 10% margin of error) for independent annotation. The first and second authors independently review release notes and supporting sources to assign integration dates. Inter-rater agreement is high (Cohen’s $k$ = 0.8524), indicating near-perfect agreement. Six disagreements are identified. These primarily arise from differing interpretations of whether certain features or naming conventions constituted evidence of generative functionality. For instance, references to “AI” in app names (e.g., “AI Song”) are not considered sufficient evidence unless accompanied by explicit indications of generative capabilities (e.g., “AI Image Generator”) or supporting descriptions. Additionally, features such as “AI effects” or “AI enhancement” are clarified as AI-assisted functionalities rather than Gen-AI because they modify content rather than generate novel content. Following the discussion, we refine the coding criteria accordingly and resolve all disagreements through consensus. The first author then applies the finalized criteria to label the remaining apps.
@@
Stage 2: Short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or fewer. Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021; Arambepola et al., 2024). This step aims to eliminate overly brief and potentially non-informative feedback.
@@
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
@@
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024). Prior research comparing explicit, implicit, and creative prompting (Giray, 2023; Ho et al., 2024; Deshmukh et al., 2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning. Consistent with these findings and with approaches such as LLM-Cure (Assi et al., 2025a), we design an explicit, structured prompt that classifies each review as either informative or non-informative. The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses. These measures ensure deterministic labeling and minimize hallucination by preventing free-form or extraneous text generation.
@@
    Iterative prompt refinement. Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin et al., 2024; He et al., 2024). In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. The final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance, is provided in Appendix B.1.